# LETS-C: Leveraging Language Embedding for Time Series Classification

### v1

@@ 4.1 Experimental Setup
Datasets and Evaluation Metrics We followed Wu et al. (2022a) and benchmarked on 10 multivariate datasets from the UEA Time Series Classification Archive (Bagnall et al., 2018), encompassing a wide range of applications, namely: EthanolConcentration (Large et al., 2018), FaceDetection (Rik Henson, 2023), Handwriting (Shokoohi-Yekta et al., 2017), Heartbeat (Liu et al., 2016), JapaneseVowels (Kudo et al., ), PEMS-SF (Cuturi, 2011), SelfRegulationSCP1 (Birbaumer et al., 1999), SelfRegulationSCP2 (Birbaumer et al., 1999), SpokenArabicDigits (Bedda & Hammami, 2010), and UWaveGestureLibrary (Liu et al., 2009). See Table 6 for their data characteristics. These datasets provide a comprehensive testing environment, featuring multivariate dimensions ranging from 3 to 963, time series lengths up to 1751, and number of classes up to 26. See Section B in the Appendix for details on each dataset.
@@
To benchmark the classifiers, we used metrics including classification accuracy and AvgWins. AvgWins is defined as the average number of times that a method outperforms other methods across benchmarked datasets, with ties also being counted towards this average. Additionally, the methods were analyzed in terms of trainable model parameters.
@@
Baselines We included 20 baseline models to ensure a comprehensive comparison. We utilize the same baselines outlined by Zhou et al. (2024); Wu et al. (2022a), namely: Classical methods: 1) Dynamic Time Warping (DTW) (Berndt & Clifford, 1994), 2) eXtreme Gradient Boosting (XGBoost) (Chen & Guestrin, 2016), and 3) RandOm Convolutional KErnel Transform (ROCKET) (Dempster et al., 2020); MLP-based methods: 4) LightTS (Zhang et al., 2022), and 5) DLinear (Zeng et al., 2023); RNN-based models: 6) Long short-term memory (LSTM) (Hochreiter & Schmidhuber, 1997), 7) LSTNet (Lai et al., 2018), and 8) LSSL (Gu et al., 2021); CNN-based models: 9) Temporal Convolutional Network (TCN) (Franceschi et al., 2019), and 10) TimesNet (Wu et al., 2022a); Transformer-based models: 11) Transformer (Vaswani et al., 2017), 12) Reformer (Kitaev et al., 2020), 13) Informer (Zhou et al., 2021), 14) Pyraformer (Liu et al., 2021), 15) Autoformer (Wu et al., 2021), 16) Stationformer (Liu et al., 2022), 17) FEDformer (Zhou et al., 2022), 18) ETSformer (Woo et al., 2022), 19) Flowformer (Wu et al., 2022b); and LLM-based model: 20) OneFitsAll (Zhou et al., 2024). Note that some of these methods were originally developed for forecasting but were later adapted for classification tasks.
Implementation Details The experiments were conducted on a Linux machine equipped with 8 NVIDIA T4 Tensor Core GPUs, each with 16GB of memory. We utilized the PyTorch v1.0.0 deep learning platform running on Python 3.6 for all models. All configurations employed the RAdam optimizer Liu et al. (2019a), with its default hyperparameters settings $(\beta_{1},\beta_{2})=(0.9,0.999)$. For tokenization, we found that maintaining a precision of one decimal place optimizes performance (see Appendix Section F). Further, we explored LETS-C’s performance across various embedding types, other than text-embedding-3-large in Section 4.3. Exploratory hyperparameter optimization was conducted, revealing that 1-4 1D convolutional layers and 1-3 linear layers are optimal for the performance of LETS-C across all datasets.
4.2 Main Results
Comparison to State-of-the-art Table 1 presents a comparative analysis of our proposed LETS-C approach against 20 baseline models across 10 benchmark datasets as introduced above.
Table 1: Comparison of classification accuracy (%) and AvgWins (%). Red: Best, Blue: Second best.
@@ 4.2 Main Results
Classical methods DTW Berndt & Clifford (1994) 32.3 52.9 28.6 71.7 94.9 71.1 77.7 53.9 96.3 90.3 66.97 0%
XGBoost Chen & Guestrin (2016) 43.7 63.3 15.8 73.2 86.5 98.3 84.6 48.9 69.6 75.9 65.98 10%
@@
RNN LSTM Hochreiter & Schmidhuber (1997) 32.3 57.7 15.2 72.2 79.7 39.9 68.9 46.6 31.9 41.2 48.56 0%
@@
Transformers Transformer Vaswani et al. (2017) 32.7 67.3 32 76.1 98.7 82.1 92.2 53.9 98.4 85.6 71.9 0%
@@
We observe that LETS-C consistently demonstrates robust performance across all datasets, achieving the highest average accuracy of 76.16% and AvgWins of 40%, compared to 20 benchmark models. This includes the most recent SOTA model OneFitsAll (accuracy: 73.97%, AvgWins: 20%) and an older SOTA TimesNet (accuracy: 73.57%, AvgWins: 0%). Notably, LETS-C surpasses OneFitsAll on six out of ten datasets by a significant margin. LETS-C is particularly effective on challenging datasets like PEMS-SF and EthanolConcentration (EC). PEMS-SF is characterized by exceptionally high dimensionality with 963 features, and EC contains an extremely long time series at length of 1751, compared to other datasets in the benchmark. These results showcase the competitive edge of LETS-C against the previous SOTA methods (OneFitsALL), thus establishing a new benchmark for time series classification.
Computational Cost Analysis Next, we aim to assess how well the LETS-C approach balances performance with computational efficiency, which is crucial for usage in resource-constrained environments. Table 2 provides a detailed analysis of the trainable parameters associated with LETS-C compared to the previous SOTA model, OneFitsAll.
Table 2: Comparison of per batch trainable parameters (millions) for LETS-C vs. OneFitsAll. The Delta (%) $=100\times\frac{\text{Cost with LETS-C{}}}{\text{Cost with OneFitsAll}}$, quantifying the efficiency of LETS-C relative to OneFitsAll.
@@
Delta (%) 19.89 0.16 8.89 2.28 11.19 5.51 30.83 32.06 7.77 26.21 14.48
@@
Our method achieved higher performance with only 14.48% of the trainable model parameters on average, compared to OneFitsAll. Despite the advantage of OneFitsAll over other leading models like TimesNet and FEDformer on its reduced parameter count, OneFitsAll still requires much more computational cost than our approach, especially for training. Further experiments on training and inference times are detailed in Table 7 in the Appendix. We show that LETS-C offers a lightweight approach to the time series classification task while achieving the SOTA performance.
4.3 Additional Analysis
Ablation Study To empirically assess the advantages of fusing both text embeddings and time series data, as opposed to variants that only leverage either the text embedding or time series itself, we conducted an ablation study.
[…]
LETS-C (Ours) 52.9 68.9 23.8 78 99.2 93.1 93.2 62.8 99.2 90.6 76.17 60%
   embedding only 38 59.4 20.1 80 99.2 92.5 88.7 63.9 99.2 88.4 72.94 40%
As shown in Table 3, we observe that the combination of both embeddings and time series achieves the highest average accuracy, at 76.17%, and the greatest number of AvgWins, compared to the ablated versions. These demonstrate the substantial performance gains from fusing embeddings with time series data, essential for optimal model accuracy.
Alternative Methods for Fusing Time Series with Embeddings We explored two additional methods for fusing time series and embeddings beyond simple addition. The first method involves a Fusion network that processes both embeddings and time series data through convolutional and dense layers, subsequently merging the features from both sources into a final dense network. The second method employs Concatenation, where the time series and embeddings are concatenated and processed through a lightweight classification head. Table 4 presents the classification accuracy and trainable model parameters for these variations.
Table 4: Comparison of classification accuracy (%) and trainable model parameters (millions) for alternative methods for fusing time series with embeddings. Higher AvgWins and averages signify superior performance, while lower averages suggest greater computational efficiency. Red: Best performance.
[…]
Concatenation 43 65.1 22.5 79 98.9 93.1 93.9 58.9 99 88.8 74.22 30%
[…]
Fusion Network 0.19 0.35 0.33 0.28 0.16 5.54 0.37 0.27 0.23 0.04 0.78
Concatenation 0.42 0.009 0.26 0.17 0.11 0.28 0.23 0.39 0.09 0.46 0.24
We observe that the addition approach in the LETS-C architecture achieves the highest average classification accuracy (76.11%) compared to the fusion network (73.40%) and concatenation (74.22%). It also records the best AvgWins at 70%. Further, the number of parameters increases with both the concatenation and fusion network approaches, resulting in additional complexity compared to the addition method. As our goal was to develop a lightweight model, we opted for the addition approach due to its simpler parameter structure.
@@ 4.3 Additional Analysis
Assessing Text Embeddings with Cosine Similarity To analyze text embeddings extracted from time series data and assess their effectiveness in distinguishing between and within data classes, we calculate the average cosine similarities for all pairs of time series within each class and similarly for between classes. Additionally, we average similarities from multiple features to effectively handle multivariate time series. This approach helps us quantify how closely time series in the embedding space are related, both within individual classes and between distinct groups. Figure 2 (left) visualizes these relationships through heatmaps. Each matrix entry is scaled using min-max normalization to range from 0 to 1, where warmer colors in the heatmap represent higher similarities and darker shades indicate lower similarities. Diagonal entries show within-class similarities, highlighting intra-class cohesion, while off-diagonal entries reveal between-class relationships.
@@
Figure 2: Left: Heatmaps illustrating within-class and between-class cosine similarities of text embeddings derived from the training time series data in JapaneseVowels (far left) with 9 classes and SpokenArabicDigits (middle) with 10 classes. Right: Trade-off between the percentage of accuracy retention and model parameter retention relative to LETS-C’s optimal values across all datasets. The optimal LETS-C accuracy (%) and parameters (millions) for each dataset are detailed in the legend.
We observe that within-class similarity consistently exceeds across-class similarity, thereby validating the hypothesis that text embeddings effectively retain and convey significant information from the underlying time series data. This confirms their utility in discerning inherent patterns within the data.
Embedding Type To explore the generalization capabilities of our approach with various embeddings beyond OpenAI’s text-embedding-3-large, we evaluated the performance of LETS-C using three distinct embedding types. These were selected based on their high rankings in the MTEB Muennighoff et al. (2022); MTEB (2024) overall leaderboard performance, varying in embedding dimensions, maximum token lengths, and model sizes. The selected embeddings and their characteristics, along with text-embedding-3-large, are detailed in Table 9, and Section G in the Appendix. Table 5 presents detailed performance metrics (accuracy) and computational cost (trainable parameters) for these various embedding models in the LETS-C framework.
We observe that LETS-C consistently outperforms current baselines in terms of average classification accuracy and AvgWins, while utilizing only a fraction of the trainable model parameters across all explored embedding types, which vary in dimensionality and model size. Specifically, text-embedding-3-large achieves an average accuracy of 76.17%, while e5-mistral-7b-instruct, gte-large-en-v1.5, and nomic-embed-text-v1 achieve accuracies of 74.89%, 76.02%, and 75.12% respectively. These figures not only surpass the previous SOTA accuracy of 73.97% but also require significantly fewer trainable model parameters—just 14.48%, 15.62%, 9.24%, and 5.31% compared to OneFitsAll. Although preliminary, we observe some patterns between the computational cost of LETS-C with an embedding model and the embedding dimensionality: e5-mistral-7b-instruct requires the highest average trainable parameters (0.25M) for the largest embedding dimensionality of 4096. This is followed by text-embedding-3-large, which has an embedding dimensionality of 3072 and needs 0.22M trainable parameters, then gte-large-en-v1.5 with a dimensionality of 1024 requiring 0.19M, and lastly nomic-embed-text-v1 with 768 dimensions needing 0.09M trainable parameters. Consequently, our approach generalizes across diverse embedding types, demonstrates superior performance, and offers a more lightweight solution, outperforming baselines across a broad range of language embedding models. Additionally, we examined the impact of time series precision on tokenization and its effect on LETS-C’s performance, detailed in Appendix Section F.
@@
Table 5: Comparison of LETS-C using various embedding models against OneFitsAll. Performances that surpass OneFitsAll are in bold with the best in Red. AvgWins scores above 50% denote consistent superiority, calculated as 1 for outperformance relative to OneFitsAll and 0 otherwise. Delta (%) $=100\times\frac{\text{Cost with each embedding model}}{\text{Cost with OneFitsAll}}$ measures computational efficiency relative to OneFitsAll.
@@
Computational Cost (Trainable Parameters (M)) $\downarrow$
@@
           Delta % 2.21 1.31 20.99 3.58 3.99 1.85 3.02 10.03 3.34 2.78 5.31
Trade-offs: Model Accuracy vs. Parameter Complexity Next, we aim to examine the trade-offs between model accuracy and parameter complexity in our approach. We explore the relationship between retained accuracy and parameter reduction across all datasets, aiming to demonstrate minimal loss in accuracy despite significant reductions in model parameters. Figure 2 (right) illustrates the trade-off between the percentage of accuracy retention and the percentage of model parameter retention, as compared to the optimal accuracy and parameters of our LETS-C model across all datasets (see Tables 1, and 7). For more detailed results, see Table 10 in the Appendix. This comparison quantifies the accuracy and computational efficiency of the reduced model relative to our best-performing model.
Figure 2 (right) underscores the efficiency of our LETS-C approach and demonstrates its effectiveness in retaining high accuracy with significantly reduced parameters across all datasets. The trade-off between model accuracy and parameter complexity is data-dependent, but generally, a substantial reduction in parameters leads to only a slight decrease in accuracy. We closely explore this trade-off in three specific datasets, namely, Heartbeat, PEMS-SF, and SpokenArabicDigits, to understand how these trade-offs manifest in different contexts, as detailed in Appendix Section H.
These examples illustrate that while fewer parameters generally lead to lower accuracy, the decrement is manageable, making LETS-C highly suitable for deployment in resource-constrained environments or applications requiring rapid processing with minimal computational overhead.

### v2

@@ 4 Experimental Protocol and Details
We evaluated LETS-C against a well-established benchmark for multivariate time series classification, commonly adopted for comparison in various studies Zhou et al. (2024); Li et al. (2024b); Wu et al. (2022a); Zerveas et al. (2021). This benchmark, selected from the UEA Archive (Bagnall et al., 2018), is purposefully curated to cover challenging and diverse domains, including the following datasets: EthanolConcentration (Large et al., 2018), FaceDetection (Rik Henson, 2023), Handwriting (Shokoohi-Yekta et al., 2017), Heartbeat (Liu et al., 2016), JapaneseVowels (Kudo et al., 1999), PEMS-SF (Cuturi, 2011), SelfRegulationSCP1 (Birbaumer et al., 1999), SelfRegulationSCP2 (Birbaumer et al., 1999), SpokenArabicDigits (Bedda and Hammami, 2010), and UWaveGestureLibrary (Liu et al., 2009). This commonly adopted benchmark offers a comprehensive testing environment, with multivariate dimensions ranging from 3 to 963, time series lengths up to 1751, and up to 26 classes. See Appendix B for details on each dataset.
@@
To assess the classifiers, we used metrics including classification accuracy and AvgWins. AvgWins is defined as the average number of times that a method outperforms other methods across benchmarked datasets, with ties also being counted towards this average. Additionally, we analyzed models’ computational efficiency in terms of trainable model parameters.
Baselines
@@
We included 27 baseline models to ensure a comprehensive comparison. We utilize the same supervised learning baselines outlined by Zhou et al. (2024); Wu et al. (2022a), namely: Classical methods: 1) Dynamic Time Warping (DTW) (Berndt and Clifford, 1994), 2) eXtreme Gradient Boosting (XGBoost) (Chen and Guestrin, 2016), and 3) RandOm Convolutional KErnel Transform (ROCKET) (Dempster et al., 2020); MLP-based methods: 4) LightTS (Zhang et al., 2022a), and 5) DLinear (Zeng et al., 2023); RNN-based models: 6) Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997), 7) Long- and Short-term Time-series Network (LSTNet) (Lai et al., 2018), and 8) Linear State Space Layer (LSSL) (Gu et al., 2021); CNN-based models: 9) Temporal Convolutional Network (TCN) (Franceschi et al., 2019), and 10) TimesNet (Wu et al., 2022a); Transformer-based models: 11) Transformer (Vaswani et al., 2017), 12) Reformer (Kitaev et al., 2020), 13) Informer (Zhou et al., 2021), 14) Pyraformer (Liu et al., 2021), 15) Autoformer (Wu et al., 2021), 16) Non-stationary Transformer (Liu et al., 2022), 17) FEDformer (Zhou et al., 2022), 18) ETSformer (Woo et al., 2022), 19) Flowformer (Wu et al., 2022b), 20) Patch Time Series Transformer (PatchTST) Nie et al. (2023); and LLM-based model: 21) OneFitsAll (Zhou et al., 2024). For details, see Appendix C. Note that some of these methods were adapted from forecasting to classification tasks, without altering the core design of the models (Appendix D). Additionally, we included the unsupervised representation learning baselines outlined by Goswami et al. (2024), namely: CNN-based methods: 22) T-Loss Franceschi et al. (2019), 23) Temporal Neighborhood Coding (TNC) Tonekaboni et al. (2021), 24) TS2Vec Yue et al. (2022); Transformer-based models: 25) Time Series representation learning framework via Temporal and Contextual Contrasting (TS-TCC) Eldele et al. (2021), 26) Time Series Transformer (TST) Zerveas et al. (2021), and 27) MOMENT Goswami et al. (2024). For details, refer to Appendices E (unsupervised baselines) and F (adapting unsupervised models for classification). Reproduction details for baselines and LETS-C implementation details are in Appendices G and H, respectively.
5 Results and Analysis
5.1 Performance and Efficiency
Comparison to State-of-the-art
Table 1 and Figure 3 present a comparative analysis of LETS-C against 27 baseline models on the benchmark introduced above [^1] To maintain consistency with prior works, we use the same settings as in TimesNet and OneFitsAll, where the UEA benchmark includes only training and testing sets (with no validation set). Consequently, the reported performance reflects the model’s upper bound measurement. We collect supervised baseline model results from Zhou et al. (2024) and Wu et al. (2022a), except for PatchTST, which is based on our own reproduction. Additionally, we gather unsupervised representation learning baseline results from Goswami et al. (2024) and Yue et al. (2022)..
Table 1: Comparison of classification accuracy (%) and AvgWins (%). Red: Best, Blue: Second best. Abbreviations: EC: Ethanol Concentration, FD: Face Detection, HW: Handwriting, HB: Heartbeat, JV: Japanese Vowels, SCP1: Self-Regulation SCP1, SCP2: Self-Regulation SCP2, SAD: Spoken Arabic Digits, UW: UWave Gesture Library.
@@ 5.1 Performance and Efficiency
Supervised Learning Methods
Classical methods DTW Berndt and Clifford (1994) 32.3 52.9 28.6 71.7 94.9 71.1 77.7 53.9 96.3 90.3 66.97 0%
XGBoost Chen and Guestrin (2016) 43.7 63.3 15.8 73.2 86.5 98.3 84.6 48.9 69.6 75.9 65.98 10%
@@
RNN LSTM Hochreiter and Schmidhuber (1997) 32.3 57.7 15.2 72.2 79.7 39.9 68.9 46.6 31.9 41.2 48.56 0%
@@
Transformer Transformer Vaswani et al. (2017) 32.7 67.3 32 76.1 98.7 82.1 92.2 53.9 98.4 85.6 71.9 0%
@@
Unsupervised Representation Learning Methods
CNN T-Loss Franceschi et al. (2019) 20.5 51.3 45.1 74.1 98.9 67.6 84.3 53.9 90.5 87.5 67.37 0%
TNC Tonekaboni et al. (2021) 29.7 53.6 24.9 74.6 97.8 69.9 79.9 55.0 93.4 75.9 65.47 0%
TS2Vec Yue et al. (2022) 30.8 50.1 51.5 68.3 98.4 68.2 81.2 57.8 98.8 90.6 69.57 0%
Transformer TS-TCC Eldele et al. (2021) 28.5 54.4 49.8 75.1 93.0 73.4 82.3 53.3 97.0 75.3 68.21 0%
TST Zerveas et al. (2021) 26.2 53.4 22.5 74.6 97.8 74.0 75.4 55.0 92.3 57.5 62.87 0%
MOMENT Goswami et al. (2024) 35.7 63.3 30.8 72.2 71.6 89.6 84.0 47.8 98.1 90.9 68.4 0%
@@
Table 2: Comparison of trainable parameters (millions) for LETS-C vs. OneFitsAll. The Ratio (%) $=100\times\frac{\text{\# of trainable parameters in {LETS-C}{}}}{\text{\# of trainable parameters in OneFitsAll}}$, quantifying the efficiency of LETS-C relative to OneFitsAll.
@@
Ratio (%) 19.89 0.16 8.89 2.28 11.19 5.51 30.83 32.06 7.77 26.21 14.48
We observe that LETS-C consistently demonstrates robust performance across all datasets, achieving the highest average accuracy of 76.16% and AvgWins of 40%, compared to 27 benchmark models. This includes the most recent SOTA model OneFitsAll (accuracy: 73.97%, AvgWins: 20%) and an older SOTA TimesNet (accuracy: 73.57%, AvgWins: 0%). Notably, LETS-C surpasses OneFitsAll on six out of ten datasets by a significant margin. LETS-C is particularly effective on challenging datasets like PEMS-SF and EthanolConcentration (EC). PEMS-SF is characterized by exceptionally high dimensionality with 963 features, and EC contains an extremely long time series at length of 1751. These results showcase the competitive edge of LETS-C against the previous SOTA methods, thus establishing a new benchmark for time series classification.
Computational Efficiency
@@
Next, we aim to assess how well LETS-C balances performance with computational efficiency, which is crucial for usage in resource-constrained environments. Table 2 provides a detailed analysis of the trainable parameters associated with LETS-C compared to the previous SOTA model, OneFitsAll. Our method achieved higher performance with only 14.48% of the trainable model parameters on average, compared to OneFitsAll. Despite the advantage of OneFitsAll over other leading models like TimesNet and FEDformer on its reduced parameter count, OneFitsAll still requires much more trainable parameters than our approach. We show that LETS-C offers a lightweight approach to time series classification while achieving the SOTA accuracy. It’s important to note that in LETS-C, the text embedding computation is a one-time operation, unlike models like OneFitsAll, which continue to incur computational costs while fine-tuning partially frozen LLMs.
5.2 Effectiveness of LETS-C
Built-in Discriminative Power of Text Embeddings on Time Series
@@ 5.2 Effectiveness of LETS-C
To analyze text embeddings extracted from time series and assess their effectiveness, we compared embeddings of time series from the same and different classes. This study revealed the built-in power of text embeddings to readily distinguish time series from different classes. Specifically, we computed the average cosine similarity for time series pairs within the same class and across different classes. We average similarities from multiple channels to handle multivariate time series effectively. This approach helps us quantify how closely time series in the embedding space are related, both within each class (intra-class) and across distinct classes (inter-class). Figure 2 (left) visualizes these embedding similarities through heatmaps. The heatmaps are scaled using min-max normalization, where warmer colors in the heatmap represent higher similarities and cooler shades indicate lower similarities. Diagonal entries show intra-class similarities, while off-diagonal entries reveal inter-class similarities.
@@
Figure 2: Left: Heatmaps illustrating within- and between-class cosine similarities of text embeddings derived from the training time series in JapaneseVowels (far left) with 9 classes and SpokenArabicDigits (middle) with 10 classes. Right: Trade-off between the percentage of accuracy retention and model parameter retention relative to LETS-C’s optimal values. The optimal LETS-C accuracy (%) and parameters (M) for each dataset are in the legend.
@@
Table 3: Comparison of LETS-C with various embedding models against OneFitsAll. Performances surpassing OneFitsAll are in bold, with the best in Red. AvgWins scores above 50% indicate consistent superiority, calculated as 1 for outperforming OneFitsAll and 0 otherwise. Ratio (%) $=100\times\frac{\text{\# of trainable parameters in {LETS-C}{}}}{\text{\# of trainable parameters in OneFitsAll}}$ measures computational efficiency relative to OneFitsAll.
@@
Trainable Parameters (M) $\downarrow$
@@
             Ratio % 2.21 1.31 20.99 3.58 3.99 1.85 3.02 10.03 3.34 2.78 5.31
We observe that intra-class similarity consistently exceeds inter-class similarity, demonstrating that text embeddings effectively retain and convey significant information from the underlying time series.
Generalization Across Various Text Embedding Models
To assess the generalization of our approach across various text embedding models beyond text-embedding-3-large, we evaluate LETS-C using three additional embeddings: e5-mistral-7b-instruct Wang et al. (2024), gte-large-en-v1.5 Li et al. (2023), and nomic-embed-text-v1 Nussbaum et al. (2024). These models were selected for their high rankings in the MTEB overall leaderboard and their variations in embedding dimensions, maximum token lengths, and model sizes. Details on embedding models are in Appendix N. Table 3 presents detailed accuracy metrics and trainable parameters for these various embedding models in the LETS-C framework. We observe that LETS-C consistently outperforms current baselines in terms of average classification accuracy and AvgWins, while utilizing only a fraction of the trainable model parameters across all explored text embedding models Specifically, text-embedding-3-large achieves an average accuracy of 76.17%, while e5-mistral-7b-instruct, gte-large-en-v1.5, & nomic-embed-text-v1 achieve accuracies of 74.89%, 76.02%, and 75.12% respectively. These results not only surpass the previous SOTA accuracy of 73.97% but also require significantly fewer trainable model parameters—just 14.48%, 15.62%, 9.24%, and 5.31% compared to OneFitsAll. Among these text embedding variations, higher text embedding dimensionality leads to more trainable parameters on average. e5-mistral-7b-instruct requires the most trainable parameters (0.25M) due to its 4096 dimensions, followed by text-embedding-3-large (3072D, 0.22M), gte-large-en-v1.5 (1024D, 0.19M), and nomic-embed-text-v1 (768D, 0.09M). Consequently, our approach generalizes across diverse text embedding models, demonstrates superior performance, with the benefit of being lightweight.
Optimizing Model Size with Minimal Accuracy Loss
Next, we examine the trade-offs between model accuracy and model size in our approach. To vary model size, we adjust the number of linear and convolution layers in the classification head, ranging from 1 to 5 layers each, creating model variants of different sizes. Across all datasets, we find that significant reductions in model parameters result in only a minimal loss of accuracy. Figure 2 (right) illustrates this trade-off, showing accuracy and parameter retention relative to the optimal performance of LETS-C (see Tables 1 and 2 for reference optimal values). The results highlight the efficiency of LETS-C, demonstrating its ability to maintain high accuracy with significantly fewer parameters across all datasets. While the trade-off between model size and accuracy is data-dependent, in general, substantial parameter reductions lead to only slight accuracy decreases. For detailed results, see Appendix O and Table 9. This analysis confirms that while reducing parameters can lower accuracy, the impact remains manageable, making LETS-C suitable for applications requiring efficient inference.
5.3 Additional Analysis
Ablation Study
To empirically evaluate the benefits of fusing text embeddings with time series over variants using only one modality, we conduct an ablation study. As shown in Appendix P and Table 10, we observe that the combination of both embeddings and time series achieves the highest average accuracy, at 76.17%, and the best number of AvgWins, compared to the ablated versions. These demonstrate the significant performance gains from fusing embeddings with time series data, which are essential for optimal model accuracy.
Alternative Methods for Fusing Time Series with Embeddings
We explore two additional methods for fusing time series and embeddings beyond simple addition. The first method involves a Fusion network that first processes embeddings and time series data through convolutional and dense layers in two separate branches, then merges the features from both branches into a final dense network. The second method employs Concatenation, where the time series and embeddings are concatenated and processed through a lightweight classification head. Despite cross-attention being another alternative for fusing different modalities, we didn’t include it in this study due to the computational complexity it adds to the model. Appendix Q and Table 11 present details on classification accuracy and trainable model parameters for these variations. We observe that the addition approach in the LETS-C architecture achieves the highest average classification accuracy (76.11%) compared to the fusion network (73.40%) and concatenation (74.22%). It also records the best AvgWins at 70%. Further, the number of parameters increases with both the concatenation and fusion network approaches, resulting in additional complexity compared to the addition method. As our goal was to develop a lightweight model, we opted for the addition approach due to its simpler parameter structure.